A vaccine requires an emulsifier called polysorbate 80 to stabilize aqueous formulations of medications. It has a density of 1.064 g/mL. Express this density in: (a) kg/m3 and (b) lb/ft3


1
Expert's answer
2021-08-26T14:48:28-0400

a)


"1.064\frac{g}{mL}=1064\frac{kg}{m^3}"

b)


"1.064\frac{g}{mL}=(1.064)62.42797=66.42\frac{lb}{ft^3}"